OpenStudy (anonymous):

Find all the real cube roots of -0.000125 Answer choices below Will Medal

OpenStudy (anonymous):

doesn't anyone own a calculator ?

OpenStudy (anonymous):

I dont know how to do it though?

OpenStudy (anonymous):

I do own a calculator

OpenStudy (anonymous):

Which one is it?

OpenStudy (anonymous):

you can do it without a calculator as well

OpenStudy (anonymous):

\[-.000125=-\frac{125}{1,000,000}\]

OpenStudy (anonymous):

the cubed root of \(125\) is \(5\) because \(5^3=125\) and the cubed root of \(1,000,000\) is \(100\)

OpenStudy (anonymous):

answer \[-\frac{5}{100}=-.05\]